# ImpliRet: Benchmarking the Implicit Fact Retrieval Challenge

### v1

@@ Abstract
Retrieval systems are central to many NLP pipelines, but often rely on surface-level cues such as keyword overlap and lexical semantic similarity. To evaluate retrieval beyond these shallow signals, recent benchmarks introduce reasoning-heavy queries; however, they primarily shift the burden to query-side processing techniques – like prompting or multi-hop retrieval – that can help resolve complexity. In contrast, we present ImpliRet, a benchmark that shifts the reasoning challenge to document-side processing: The queries are simple, but relevance depends on facts stated implicitly in documents through temporal (e.g., resolving “two days ago”), arithmetic, and world knowledge relationships. We evaluate a range of sparse and dense retrievers, all of which struggle in this setting: the best nDCG@10 is only 15.07%. We also test whether long-context models can overcome this limitation. But even with a short context of only ten documents, including the positive document, GPT-4.1 scores only 35.06%, showing that document-side reasoning remains a challenge. Our codes are available at github.com/ZeinabTaghavi/IMPLIRET.††* Equal Contribution.
@@ 1 Introduction
In contrast, we present ImpliRet, a benchmark that shifts reasoning to document-side processing: the queries are simple, but relevance depends on facts stated implicitly within the documents, spanning arithmetic, temporal, and world knowledge relationships that require inference to uncover. Figure 1 gives an example: the correct document requires resolving a reference to a date that is implicit, i.e., not stated directly. An effective retrieval system must infer such implicit facts from the document content, ideally as part of the indexing process, in order to retrieve the correct result at query time. Yet current retrieval methods fail to capture the implicit signals needed for accurate retrieval. We evaluate sparse and dense approaches, including BM25 Robertson and Zaragoza (2009), ColBERT Santhanam et al. (2022), and Dragon+ Lin et al. (2023), and observe consistently poor performance: the best nDCG@10 is only 15.07% across our benchmark. To test whether long-context capabilities could mitigate the problem, we evaluate models in a setting where the positive document is included among several distractors. While GPT-4.1 answers correctly when given only the positive document, its performance drops sharply even with just ten documents in-context, achieving a ROUGE-1 recall of 35.06%. Our dataset ImpliRet introduces a new setting that requires document-side reasoning for retrieval rather than query-side reasoning. ImpliRet presents challenges for both retrieval and long-context processing, highlighting the need for models that can reason over implicit information embedded in large corpora.
@@ 2.1 Generating Tuple Sets
An arithmetic relation requires simple numerical reasoning. For instance, the query “Which bag costs $1,600?” can be answered by “The Prada bag costs $2,000, the Gucci bag is 20% cheaper,” since $2,000 × 0.8 = $1600. Here, the model must identify the reference price, interpret the relative statement (“20% cheaper”), and perform the corresponding computation to infer the answer. Therefore, each tuple in the implicit tuple set takes the form , where is the base price, is the relative multiplier, indicates the direction of the change, and is the queried price (e.g., ). We apply constraints to ensure that queried prices are unique, realistic, and well-distributed across the tuple set. Tuples are generated using a sampling algorithm that selects base prices and checks constraint satisfaction, backtracking as needed until valid tuples are found (where is the target number of documents indicated as “Docs” in Table 1). Full constraint details and sampling logic are provided in Appendix 2.1.
@@ 2.2 Document-Query Pairs
In both styles, i.e., in each conversation and post, every message includes a timestamp and speaker name (see Figure 1). In both styles, each example is produced via a three-step pipeline: (1) Entity binding: We assign entities (e.g., names, items, activities) to each tuple to create a plausible scenario and define the query target; (2) Document generation: We prompt an LLM to generate a chat or forum passage that embeds the entity and the implicit part of the tuple, without stating the explicit fact; (3) Verification: a second model attempts to extract the original tuple; we retain only examples where the intended fact is fully recoverable. This pipeline is supported by auxiliary lexical resources, including random names, brand-item pairs, and activity lists, as well as per-reasoning category prompt templates. We use LLAMA 3.3-70B Meta (2024) to synthesize the documents for each tuple.111Details such as prompts and query templates are available in Appendix A.4. Table 1 presents ImpliRet statistics222The tokens are counted using GPT-2 tokenizer Radford et al. (2019)..
@@ 3 Experiments
Contriever | 16.15 | 13.84 | 12.59 | 14.19 |
Dragon+ | 17.15 | 14.61 | 12.53 | 14.76 |
ReasonIR | 19.53 | 10.74 | 14.93 | 15.07 |
@@ 4 Results
The nDCG@10 results across all reasoning categories are presented in Table 2. The highest average score, 15.07 (achieved by ReasonIR, a recent 8B-parameter LLM), shows the difficulty retrieval models face when reasoning over implicit facts in documents. More efficient baselines such as Contriever, Dragon+, and BM25 perform substantially worse; notably, BM25 reaches just 12.13 due to its reliance on surface-level lexical overlap.
Performance varies across reasoning types: the World Knowledge category exhibits the largest performance spread (14.10 vs. 19.53), while it is narrowest for Arithmetic (10.74 vs. 14.61). Discourse style also plays a role: ReasonIR scores 20.58 on multi-speaker examples compared to 9.56 on uni-speaker ones, suggesting that stylistic structure affects retrieval difficulty.333Full results per category and style in Appendix B.
@@
We evaluate three settings: (positive only), (positive plus nine negatives), and a full-pool setting where all documents from the pool are provided as context. The model receives the query along with the sequence of documents and must generate an answer. We evaluate two reader models: LLAMA 3.3 70B and GPT-4.1.444Checkpoint: gpt-4.1-2025-04-14 In Table 3, we report the average ROUGE-1 recall555R-1 Rec. scores to measure the overlap between the generated output and the positive answer Lin (2004). When given only the positive document (), the two models achieve average ROUGE-1 Recall of 81.92 and 88.05. This suggests that the query itself is straightforward to answer once the relevant document is isolated. This also means that an LLM can solve the task if a high-performing retriever (which would retrieve the relevant document at rank 1) is available. However, as increases (even with the positive included), performance declines, showing that LLMs struggle to focus on the correct evidence amid structurally similar negatives. This supports prior findings on long-context limitations and highlights the need for retrieving a small, focused set of documents rather than increasing context size Kuratov et al. (2024); Modarressi et al. (2025).
@@ 5 Conclusion
We introduce ImpliRet, a benchmark for evaluating retrieval models when relevance depends on document-side reasoning on implicit facts. Unlike prior datasets that emphasize complex queries, ImpliRet shifts the reasoning burden to the documents. It covers three reasoning types – world knowledge, arithmetic, and temporal – and two discourse styles. Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 15.07. Even with GPT-4.1, given ten documents including the gold, performance peaks at just 35.06%. These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
@@ A.2 World-Knowledge Reasoning
As described in Section 2, we gather landmark-country pairs under three constraints: (i) each landmark must refer to a globally unique location, avoiding names that could correspond to multiple places; (ii) the landmark name must not include lexical, semantic, or language-specific cues that reveal its country, avoiding surface-form shortcuts; and (iii) landmarks must refer to specific, recognizable sites rather than generic ones. To do so, we first assemble a seed list of unique landmark-country pairs via five steps: (i) issue a SPARQL query to Wikidata to retrieve every entity whose instance-of (P31) chain includes exactly one of the high-level place classes, museum (Q33506), university (Q3918), church building (Q16970), venue (Q17350442), or landmark (the superclass of Q17350442), and that is linked to exactly one sovereign state via the country (P17) property; (ii) for each hit, extract the English labels of its enclosing administrative region (P131), city (P131 restricted to Q515), generic location (P276), and street (P669), yielding up to five concentric location strings; (iii) discard entities with missing, machine-generated, or multi-country labels, then drop any whose name tokens overlap these location strings; embed each remaining landmark-country pair with the 768-dimensional Contriever encoder and retain only those with cosine similarity below 0.25; (iv) pass the remainders to a 70B-parameter Llama-3.3 classifier that flags and removes generic names (e.g., “Downtown Club”) or labels leaking their country, using exponential back-off retries until accepted; and (v) submit each accepted label to GPT-4o in web-search mode666Checkpoint: gpt-4o-search-preview-2025-03-11, prompting it to return the place’s country in a dictionary format, and keep only those for which the model returns exactly one location. This process yields a balanced pool of 100 unique landmark-country pairs, ensuring we sample across different countries rather than multiple landmarks from the same country. Finally, for each implicit tuple set, we select a set of distinct countries and, for each , sample one landmark from that country’s filtered list . The resulting implicit tuple sets is as follows:
@@ A.4 Synthesizing the Context
Now, for each reasoning category and document style, we have a list of implicit tuple sets containing all the information needed to have a consistent dataset. Consider having the auxiliary lexical content (personal names, daily-work verbs, brand names with corresponding items, and per-category forum topics and questions needed)777Generated using GPT-o1 OpenAI (2024), to each tuple of the implicit tuple list, we assign unique entities and then each tuple, contains all the information to generating the document in natural way. The exact item required for each reasoning category and style is mentioned in Table 4. Depending on the style, we proceed as follows:
@@ Appendix B Results
Contriever | 10.85 | 8.97 | 8.18 | 9.33 |
Dragon+ | 11.80 | 9.53 | 8.30 | 9.88 |
ReasonIR | 14.67 | 7.04 | 10.14 | 10.62 |
@@
MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | MRR@ | nDCG@ | |
@@
Contriever | 9.88 | 14.80 | 11.81 | 17.51 | 8.49 | 13.17 | 9.45 | 14.51 | 7.58 | 11.91 | 8.79 | 13.27 | 8.65 | 13.29 | 10.01 | 15.09 |
Dragon+ | 10.00 | 14.92 | 13.59 | 19.38 | 9.24 | 14.10 | 9.83 | 15.11 | 7.30 | 10.93 | 9.31 | 14.13 | 8.85 | 13.32 | 10.91 | 16.21 |
ReasonIR | 8.12 | 11.50 | 21.21 | 27.57 | 3.71 | 5.66 | 10.37 | 15.81 | 7.70 | 11.51 | 12.59 | 18.35 | 6.51 | 9.56 | 14.72 | 20.58 |

### v2

@@ Abstract
Retrieval systems are central to many NLP pipelines, but often rely on surface-level cues such as keyword overlap and lexical semantic similarity. To evaluate retrieval beyond these shallow signals, recent benchmarks introduce reasoning-heavy queries; however, they primarily shift the burden to query-side processing techniques – like prompting or multi-hop retrieval – that can help resolve complexity. In contrast, we present ImpliRet, a benchmark that shifts the reasoning challenge to document-side processing: The queries are simple, but relevance depends on facts stated implicitly in documents through temporal (e.g., resolving “two days ago”), arithmetic, and world knowledge relationships. We evaluate a range of sparse and dense retrievers, all of which struggle in this setting: the best nDCG@10 is only 14.91%. We also test whether long-context models can overcome this limitation. But even with a short context of only thirty documents, including the positive document, GPT-o4-mini scores only 55.54%, showing that document-side reasoning remains a challenge. Our codes are available at github.com/ZeinabTaghavi/IMPLIRET.††* Equal Contribution.
@@ 1 Introduction
In contrast, we present ImpliRet, a benchmark that shifts reasoning to document-side processing: the queries are simple, but relevance depends on facts stated implicitly within the documents, spanning arithmetic, temporal, and world knowledge relationships that require inference to uncover. Figure 1 gives an example: the correct document requires resolving a reference to a date that is implicit, i.e., not stated directly. An effective retrieval system must infer such implicit facts from the document content, ideally as part of the indexing process, in order to retrieve the correct result at query time. Yet current retrieval methods fail to capture the implicit signals needed for accurate retrieval. We evaluate sparse and dense approaches, including BM25 Robertson and Zaragoza (2009), ColBERT Santhanam et al. (2022), and Dragon+ Lin et al. (2023), and observe consistently poor performance: the best nDCG@10 is only 14.91% across our benchmark. To test whether long-context capabilities could mitigate the problem, we evaluate models in a setting where the positive document is included among several distractors. While GPT-o4-mini answers correctly when given only the positive document, its performance drops sharply even with just thirty documents in-context, achieving a ROUGE-1 recall of 55.54%. Our dataset ImpliRet introduces a new setting that requires document-side reasoning for retrieval rather than query-side reasoning. ImpliRet presents challenges for both retrieval and long-context processing, highlighting the need for models that can reason over implicit information embedded in large corpora.
@@ 2.1 Generating Tuple Sets
An arithmetic relation requires simple numerical reasoning. For instance, the query “Which bag costs $1,600?” can be answered by “The Prada bag costs $2,000, the Gucci bag is 20% cheaper,” since $2,000 × 0.8 = $1600. Here, the model must identify the reference price, interpret the relative statement (“20% cheaper”), and perform the corresponding computation to infer the answer. Therefore, each tuple in the implicit tuple set takes the form , where is the base price, is the relative multiplier, indicates the direction of the change, and is the queried price (e.g., ). We apply constraints to ensure that queried prices are unique, realistic, and well-distributed across the tuple set. Tuples are generated using a sampling algorithm that selects base prices and checks constraint satisfaction, backtracking as needed until valid tuples are found (where is the target number of documents indicated as “Docs” in Table 1). Full constraint details and sampling logic are provided in Appendix A.1.
@@ 2.2 Document-Query Pairs
In both styles, i.e., in each conversation and post, every message includes a timestamp and speaker name (see Figure 1). In both styles, each example is produced via a three-step pipeline: (1) Entity binding: We assign entities (e.g., names, items, activities) to each tuple to create a plausible scenario and define the query target; (2) Document generation: We prompt an LLM to generate a chat or forum passage that embeds the entity and the implicit part of the tuple, without stating the explicit fact; (3) Verification: a second model attempts to extract the original tuple; we retain only examples where the intended fact is fully recoverable. This pipeline is supported by auxiliary lexical resources, including random names, brand-item pairs, and activity lists, as well as per-reasoning category prompt templates. We use Gemma-3-27B-it Team (2025) to synthesize the documents for each tuple.111Details such as prompts and query templates are available in Appendix A.4. Table 1 presents ImpliRet statistics222The tokens are counted using GPT-2 tokenizer Radford et al. (2019)..
@@ 3 Experiments
Contriever | 16.50 | 13.70 | 12.73 | 14.31 |
Dragon+ | 17.46 | 14.61 | 12.66 | 14.91 |
ReasonIR | 18.88 | 10.78 | 11.25 | 13.64 |
@@ 4 Results
The nDCG@10 results across all reasoning categories are presented in Table 2. The highest average score, 14.91 (achieved by Dragon+), shows the difficulty retrieval models face when reasoning over implicit facts in documents. More efficient baselines such as Contriever and BM25 perform substantially worse; notably, BM25 reaches just 12.24 due to its reliance on surface-level lexical overlap.
Performance varies across reasoning types: the Arithmetic category exhibits the largest performance spread (14.96 vs. 10.78), while it is narrowest for Temporal (12.83 vs. 10.98). Discourse style also plays a role: Dragon+ scores 16.45% on multi-speaker examples compared to 13.37 on uni-speaker ones, suggesting that stylistic structure affects retrieval difficulty.333Full results per category and style in Appendix B.
@@
We evaluate three settings: (positive only), (positive plus nine negatives), and (a full-pool setting where all documents from the pool are provided as context). The model receives the query along with the sequence of documents and must generate an answer. We evaluate three reader models: LLAMA 3.3 70B, GPT-4.1444Checkpoint: gpt-4.1-2025-04-14, and GPT-o4-mini555Checkpoint: o4-mini-2025-04-16. In Table 3, we report the average ROUGE-1 recall666R-1 Rec. scores to measure the overlap between the generated output and the positive answer Lin (2004). When given only the positive document (), the models achieve average ROUGE-1 Recall of 81.92, 88.05, and 92.74. This suggests that the query itself is straightforward to answer once the relevant document is isolated. This also means that an LLM can solve the task if a high-performing retriever (which would retrieve the relevant document at rank 1) is available. However, as increases (even with the positive included), performance declines, showing that LLMs struggle to focus on the correct evidence amid structurally similar negatives. This supports prior findings on long-context limitations and highlights the need for retrieving a small, focused set of documents rather than increasing context size Kuratov et al. (2024); Modarressi et al. (2025).
@@ 5 Conclusion
We introduce ImpliRet, a benchmark for evaluating retrieval models when relevance depends on document-side reasoning on implicit facts. Unlike prior datasets that emphasize complex queries, ImpliRet shifts the reasoning burden to the documents. It covers three reasoning types – world knowledge, arithmetic, and temporal – and two discourse styles. Across sparse, dense, and KG-augmented retrievers, the best nDCG@10 is only 14.91. Even with GPT-o4-mini, given thirty documents including the gold, performance peaks at just 55.54%. These results highlight the difficulty of retrieving implicit facts and the need for models that can reason beyond surface cues.
@@ A.2 World-Knowledge Reasoning
As described in Section 2, we gather landmark-country pairs under three constraints: (i) each landmark must refer to a globally unique location, avoiding names that could correspond to multiple places; (ii) the landmark name must not include lexical, semantic, or language-specific cues that reveal its country, avoiding surface-form shortcuts; and (iii) landmarks must refer to specific, recognizable sites rather than generic ones. To do so, we first assemble a seed list of unique landmark-country pairs via five steps: (i) issue a SPARQL query to Wikidata to retrieve every entity whose instance-of (P31) chain includes exactly one of the high-level place classes, museum (Q33506), university (Q3918), church building (Q16970), venue (Q17350442), or landmark (the superclass of Q17350442), and that is linked to exactly one sovereign state via the country (P17) property; (ii) for each hit, extract the English labels of its enclosing administrative region (P131), city (P131 restricted to Q515), generic location (P276), and street (P669), yielding up to five concentric location strings; (iii) discard entities with missing, machine-generated, or multi-country labels, then drop any whose name tokens overlap these location strings; embed each remaining landmark-country pair with the 768-dimensional Contriever encoder and retain only those with cosine similarity below 0.25; (iv) pass the remainders to a 70B-parameter Llama-3.3 classifier that flags and removes generic names (e.g., “Downtown Club”) or labels leaking their country, using exponential back-off retries until accepted; and (v) submit each accepted label to GPT-4o in web-search mode777Checkpoint: gpt-4o-search-preview-2025-03-11, prompting it to return the place’s country in a dictionary format, and keep only those for which the model returns exactly one location. This process yields a balanced pool of 100 unique landmark-country pairs, ensuring we sample across different countries rather than multiple landmarks from the same country. Finally, for each implicit tuple set, we select a set of distinct countries and, for each , sample one landmark from that country’s filtered list . The resulting implicit tuple sets is as follows:
@@ A.4 Synthesizing the Context
Now, for each reasoning category and document style, we have a list of implicit tuple sets containing all the information needed to have a consistent dataset. Consider having the auxiliary lexical content (personal names, daily-work verbs, brand names with corresponding items, and per-category forum topics and questions needed)888Generated using Gemma-3-27B-it Team (2025), to each tuple of the implicit tuple list, we assign unique entities and then each tuple, contains all the information to generating the document in natural way. The exact item required for each reasoning category and style is mentioned in Table 4. Depending on the style, we proceed as follows:
@@ Appendix B Results
Contriever | 11.19 | 8.84 | 8.48 | 9.50 |
Dragon+ | 11.97 | 9.47 | 8.26 | 9.90 |
ReasonIR | 14.23 | 7.13 | 7.78 | 9.71 |
@@
MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | MRR@10 | nDCG@10 | |
@@
Contriever | 9.77 | 14.77 | 12.60 | 18.23 | 8.23 | 12.87 | 9.45 | 14.52 | 8.17 | 12.22 | 8.78 | 13.25 | 8.72 | 13.29 | 10.28 | 15.33 |
Dragon+ | 9.94 | 14.85 | 14.00 | 20.07 | 9.11 | 14.11 | 9.83 | 15.12 | 7.20 | 11.16 | 9.32 | 14.16 | 8.75 | 13.37 | 11.05 | 16.45 |
ReasonIR | 7.75 | 10.54 | 20.71 | 27.21 | 3.87 | 5.68 | 10.39 | 15.89 | 2.95 | 4.17 | 12.60 | 18.33 | 4.86 | 6.80 | 14.57 | 20.48 |